OpenStudy (anonymous):

write the equation of the line which passes through (-4,2) and is parallel to y=-x+6 in slope intercept form

OpenStudy (saifoo.khan):

parallel means same slope.

OpenStudy (saifoo.khan):

2 = -4(-1) + c c = 2 -4 c = -2 therefore y = -1x -2

OpenStudy (anonymous):

what is c? and what formula did you use?

OpenStudy (saifoo.khan):

C is the y-intercept.

OpenStudy (anonymous):

what formula?

OpenStudy (anonymous):

thank you

OpenStudy (saifoo.khan):

http://www.mathsnet.net/graphs/images/y=mxplusc.gif

OpenStudy (anonymous):

i think saifoo mean y=mx+c which m is gradient and c is constant which in this equation mean y-intercept

OpenStudy (saifoo.khan):

Yeah.